OpenStudy (anonymous):

Find the solution to the differential equation dy/dx = (1+x^2 ) / (cos (y)) with initial condition, y(0) = pi/2 . Make sure to isolate y.

OpenStudy (anonymous):

\[\begin{align*}\frac{dy}{dx}&=\frac{1+x^2}{\cos y}\\ \cos y~dy&=(1+x^2)~dx \end{align*}\] Integrate both sides, you get \[\sin y=x+\frac{1}{3}x^3+C\] Plug in the given initial condition to solve for \(C\).

OpenStudy (anonymous):

Thanks A lot.

OpenStudy (anonymous):

yw